OpenStudy (mathstudent55):

What are complementary angles? The definition?

OpenStudy (anonymous):

Either of two angles whose sum is 90°.

OpenStudy (mathstudent55):

Right. Complementary angles are two angles the sum of whose measures is 90.

OpenStudy (anonymous):

here are the answer choices. a 90. b. 180. c. 19. d. 25

OpenStudy (mathstudent55):

So you know that m<M + m<V = 90 You also know that m<M = 3x - 14, and m<V = 2x + 9, so replace m<M and m<V in m<M + m<V = 90 by the expressions in x and solve for x.

OpenStudy (anonymous):

can you just give me the answer please

OpenStudy (mathstudent55):

That wouldn't be any fun for you. The whole fun in math is to solve a problem and feel great that you did it. I'll guide you step by step.

OpenStudy (mathstudent55):

m<M + m<V = 90 m<M = 3x - 14, and m<V = 2x + 9 Do the substitutions to get: 3x - 14 + 2x + 9 = 90

OpenStudy (mathstudent55):

Now combine like terms on left side

OpenStudy (mathstudent55):

5x - 5 = 90 Now add 5 to both sides. Then divide both sides by 5. What do you get for x?

OpenStudy (anonymous):

5x-5=90 +5 +5 10x=90 /5 /5 2 18 and then what?

OpenStudy (anonymous):

oh wait it supposed to be 95 isnt it?

OpenStudy (mathstudent55):

Wait, 90 + 5 = 95, not 90. Also, on the left side, you have 5x - 5. Whan you add 5 to it, you get 5x, not 10x. 5x - 5 = 90 Add 5 to both sides: 5x = 95 Now divide both sides by 5 to get x.

OpenStudy (anonymous):

thank you so much. i got 19 :)

OpenStudy (mathstudent55):

That's it! Good job! You're welcome.

OpenStudy (anonymous):

haha thanks (: